OpenStudy (anonymous):

Find the derivative of f(x) = -12x^2 + 9x at x = 6.

OpenStudy (aravindg):

Can you first find the derivative of -12x^2 + 9x?

OpenStudy (anonymous):

I'm not sure how to do this type of problem

OpenStudy (aravindg):

okay differential of sum is sum of differentials d/dx(-12x^2)=-12 * d/dx(x^2) Can you do that part?

OpenStudy (anonymous):

is there an easier way to do this? i don't understand the d/dx

OpenStudy (aravindg):

d/dx means differential with respect to x.

OpenStudy (anonymous):

i still don't understand how to do it

OpenStudy (aravindg):

Do you know the differential of x^n?

OpenStudy (anonymous):

the notes i have did it in more of a step by step simpler way, but in my notes i only have an equation that begins with x^2, not necessarily with a coefficient in front of it

OpenStudy (aravindg):

Okay the differential of x^n is nx^n-1 You could check it for x^2 d/dx(x^2)=(2)(x)^(2-1)=2x

OpenStudy (anonymous):

i think i got it, is it -135? @AravindG

OpenStudy (aravindg):

Wonderful! :) You are right

OpenStudy (anonymous):

thanks a lot!